solve using the princible of zero products
0=s(s+3)




Found 2 solutions by stanbon, AnlytcPhil:
Answer by stanbon(75887)

Since the product is zero, one of the factors must be zero.
Therefor s=0 or s+3=0
Solutions: s=0 or s=-3
Cheers,
Stan H.

Answer by AnlytcPhil(1810)

solve using the princible of zero
products

0 = s(s + 3) 


One of the factors s and/or (s + 3) must be 0 since their
product is zero

Case 1:  s = 0
Case 2:  s + 3 = 0 or s = -3

So there are two possible solutions, s = 0 and s = -3

Checking s = 0

0 = s(s + 3)
0 = 0(0 + 3)
0 = 0(3)
0 = 0

checking s = -3

0 = s(s + 3)
0 = -3(-3 + 3)
0 = -3(0)
0 = 0

Edwin
AnlytcPhil@aol.com